# Unified Gaussian Primitives for Scene Representation and Rendering

### v1

@@ 9. Conclusion
Our method has several limitations that could serve as fruitful topics for future research. Our transmittance model shares the common limitation with the model by Vicini et al. (2021) that it does not conform to certain physical constraints, such as the weak reciprocity proposed by d’Eon (2018). Developing a reciprocal formulation for general heterogeneous non-exponential transport remains an open problem. Additionally, our phase function currently does not support refraction. Rendering refraction requires tracking the change of index of refraction when a ray enters or exits a medium boundary. When a scene is entirely modeled by our volumetric Gaussian primitives, there is no clear definition of medium boundaries and mechanism to separate interior and exterior parts. It would also be desirable to further extend our phase function to support advanced effects such as subsurface scattering. While out of scope in this work, a full path-space differentiable rendering formulation for our representation, similar to that for exponential volumes (Zhang et al., 2021b), will enable more powerful inverse rendering applications. Finally, a GPU implementation of our method will likely achieve significant performance improvement over our current CPU implementation.

### v2

@@ 9. Conclusion
Our method has several limitations that could serve as fruitful topics for future research. Our transmittance model shares the common limitation with the model by Vicini et al. (2021) that it does not conform to certain physical constraints, such as the weak reciprocity proposed by d’Eon (2018). Developing a reciprocal formulation for general heterogeneous non-exponential transport remains an open problem.
Additionally, our method currently does not support refraction inside a solid. Rendering refraction requires tracking the change of index of refraction when a ray enters or exits a medium boundary. When a scene is entirely modeled by our volumetric Gaussian primitives, there is no clear definition of medium boundaries or mechanism to separate interior and exterior parts. However, translucency is possible using a thin-surface BSDF as the base BSDF of our phase function (Eq. 15). Fig. 15 shows a simple example.
Our method currently assumes that scattering in a primitive can be modeled accurately without an exitant positional distribution, as considered in subsurface scattering (Jensen et al., 2001) or shell transport (Moon et al., 2007). This is justified by the typical use case where a single primitive covers a small surface patch or a small cluster of oriented elements. A more stringent treatment is left for future applications such as level-of-detail rendering, where the impact of the exitant positional distribution could be more obvious as a primitive covers a large volume (Kerbl et al., 2024).
While out of the scope of this work, a full path-space differentiable rendering formulation for our representation, similar to that for exponential volumes (Zhang et al., 2021b), will enable more powerful inverse rendering applications. Finally, a GPU implementation of our method will likely achieve significant performance improvement on intersection testing over our current CPU implementation, as shown by recent work (Moenne-Loccoz et al., 2024).
@@ Appendix C Magnitude Remapping from 3DGS
Appendix D Sampling Exponential Transmittance Gaussians via Analog Decomposition Tracking
Kutz et al. (2017) show that by decomposing an exponential medium into components, a free-flight sample from the original medium can be obtained by taking the minimum of independent free-flight samples from the components. This technique can thus be adopted to sample the exponential variants of scenes in Fig. 12. Each is a heterogeneous exponential medium where the extinction coefficient is defined by a mixture of 3D Gaussians: $\sigma_{\mathrm{t}}(x)=\sum_{k}G_{k}(x)$. For each ray, we simply iterate all its intersecting Gaussians and sample each of them independently. Let $u\in[0,1)$ be a random number and $x_{t}=x+t\omega,t\in[t_{0},t_{1}]$ be a ray. Sampling the exponential free-flight distance from one Gaussian is straightforward by inverting Eq. 4 and Eq. 10:
(25) $u=1-\exp(-I(t_{0},t))\Leftrightarrow t=I^{-1}(-\ln(1-u);t_{0}).$
Finally, we take the nearest sample among samples from all Gaussians.
It is important to recognize the difference between this approach and our approach with linear transmittance Gaussian primitives. Such approach falls under the conventional category of specifying a volume using per-point phase function, as discussed in §4. While the minimum operator in analog decomposition tracking conceptually resembles “selecting a primitive”, it still samples an infinitesimal collision point based on the continuous free-flight PDF. Multiple scattering within the same primitive may happen in subsequent bounces. This is different from our approach where we sample an entire primitive and never explicitly account for sub-primitive multiple scattering (Fig. 2).